HW10: Problem 7 Prev Up Next (1 pt) Let V be the vector space P3z of polynomials in with degree less than 3 and W be the subspace a. Find a nonzero polynomial p(z) in w p(z) b. Find a polynomial q(z) in V\W g(z) Note: You can earn partial credit on this problem. Preview Answers Submit Answers
Answer #1

(a). Consider p(z) = 22.2-10-r + 1 .

Then, p(x)=1.(7-9x-4x^2)+1.[6x^2-(6+x)] \in span \ W , and p(x) \neq 0 .

(b) Consider q(x)=x . Then q(x) \in V as q(x) is a polynomial in V with degree less than 3 .

We show that q(x) \notin W .

To do this, assume that q(x) \in W . Then q(x) =x=a.(7-9x-4x^2)+b[6x^2-(6+x)] for some scalars a and b .

\Rightarrow x=x^2(6b-4a)+x(-9a-6b)+(7a-6b).

Comparing coefficients of equal powers on both sides of te above equation, we have

6b-4a=0 \\ 9a+6b=0 \\ 7a-6b=0.Solving the first and third equations, we get b= \frac{2}{3}a and b= \frac{7}{6}a respectively,

Therefore, \frac{2}{3}a= \frac{7}{6}a \ \Rightarrow a=0 . Then, from the firat equation, we get b=0 .

But, this gives us q(x) =x=0.(7-9x-4x^2)+0[6x^2-(6+x)] , which is not possible and therefore a contadiction to our assumtion that q(x) \in W .

Thus, we have shown that q(x) \in V and q(x) \notin W . Hence q(x) \in V\setminus W .

P.S. : Please upvote if you have found this anwer to be useful.